40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4294     Entry time: Tue Feb 15 02:13:16 2011
Author: kiwamu 
Type: Summary 
Category: ASC 
Subject: a daytime task : small signals on ETMX OL 

Rana and I found that the QPD for the optical lever at X end are showing small signals.

At this moment each of the segments exhibits approximately 200 counts when the oplev beam is centered.

These small numbers may be due to the coating of ETMX, but we are not sure.

Probably we have to increase the gain of the QPD depending on situations.

 

So a set of the tomorrow's daytime task is:

   1. check the trend data of the QPD outputs to see how much signals were there in the past.

   2. check the whitening filters to make sure if it's on or off.

   3. If it's necessary, increase the gain of the QPD to have reasonable readouts.

I am going to ask somebody to do this task.

ELOG V3.1.3-